OpenStudy (anonymous):

what's 6 (1+5y)=30y-2

OpenStudy (anonymous):

You can multiply the 6(1+5y) to make it easier for you. Your equation then becomes 6 + 30y = 30y - 2. You're solving for y, correct? If so, you need y on one side and the number on the other. Add two to both sides to cancel out the 2 on the right side. Your equation is then 8 + 30y = 30y. Wait.. you gotta subtract 30y but then you'd have 0y...| Are there multiple choice answers?

OpenStudy (shreenithaa):

6(1+5y)=30y-2 6+30y=30y-2 -30y -30y 6=-2 the solution is no solution